OpenStudy (anonymous):

Can someone help me make a trinominal and put it in ascending order and find the degree?

OpenStudy (anonymous):

A trinomial is a polynomial that consists of three nominals like \[a^3 + b^2 + c\]. The power of the highest exponent reveals the degree of the polynomial.

OpenStudy (anonymous):

so 6^3 + 2^2 + 8 is an appropriate answer?

OpenStudy (anonymous):

and the degree would be 3?

Directrix (directrix):

@ MsCrossCounty --> A trinomiall has 3 terms. You need a variable. 6 x^3 + 2 x^2 + 8 = P(x) will do BUT....

Directrix (directrix):

.. it is not in ascending order.

OpenStudy (anonymous):

How do I put it in ascending order?

OpenStudy (anonymous):

??

Directrix (directrix):

It is now in descending order which is not the way polynomials usually are wriiten. In ascending order (low to high), the terms go from the lowest degree to the highest. P(x) = 8 + 2 x^2 + 6 X^3 The degree of the polynomial is 3 because it is the greatest exponent of the group. This polynomial is known as a cubic trinomial. Do check my work and see if it makes sense.

OpenStudy (anonymous):

it makes sense, thank you so much